OpenStudy (anonymous):

X^2 + 8x = 10 Solve by completing the square? ill give medals

OpenStudy (anonymous):

\[ x^2 + 8x = 10 \\ x^2+8x + 16 = 10+16=26\\ (x+4)^2 =26 \] Can you finish it?

OpenStudy (anonymous):

uhh i dont think sooo....

OpenStudy (anonymous):

\[ (x+4)^2=26\\ x+4 = \pm \sqrt{26}\\ x= -4\pm \sqrt{26} \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

yes thank you !!